ROTATIONALLY RESOLVED ELECTRONIC SPECTRUM OF THE PARA-DIFLUOROBENZENE-NITROGEN VAN DER WAALS COMPLEX: STRUCTURE AND INTERNAL MOTION^{a}

\underline{CHEOLHWA KANG} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260; MARTIN SCHÄFER, Laboratorium für Physikalische Chemie, Eidgenössische Technische Höchschule Zürich, Zürich, ETH Hönggerberg - HCI, CH-8093 Zürich, Switzerland.

The rotationally resolved S_{1} ← S_{0} fluorescence excitation spectra of p-difluorobenzene and its N_{2} complex have been observed. While the spectrum of pDFB is pure b-type, the spectrum of the N_{2} complex is a superposition of two pure c-type spectra that result from the internal rotation of the N_{2} molecule. The analysis reveals that N_{2} molecule lies above the center of the ring plane in the S_{0} state; this distance decreases in the S_{1} state. The preferred orientation and barrier to internal rotation in the two electronic states are also different. A full discussion of the factors responsible for the differences will be given.

awork supported by NSF.